OpenStudy (moazzam07):

)A hyena can run one over four of a mile in 22.5 seconds. At this rate, which equation can be used to determine how fast a hyena runs in miles per hour? one-fourth mile over twenty-two point five seconds = 40 miles per hour twenty-two point five seconds over one-fourth mile = 40 miles per hour one-fourth mile over twenty-two point five seconds = 90 miles per hour twenty-two point five seconds over one-fourth mile = 90 miles per hour

OpenStudy (zpupster):

\[(\frac{ .25mi }{ 22.5s })(\frac{ 60s }{ 1\min })(\frac{ 60\min }{ 1hr })=?mi/hr\]

OpenStudy (moazzam07):

thanks